find "y", when "y= sin x - cos x"


1
Expert's answer
2021-02-01T05:45:05-0500

Solution:

"The\:compound\:periodicity\:of\:the\:sum\:of\:periodic\:functions\:is\:the\:least\:\\common\:multiplier\:of\:the\:periods"


"Periodicity\;of\;sin(x):\;2\pi\\Periodicity\;of\;cos(x):\;2\pi"



"\mathrm{Domain\:of\:}\:\sin \left(x\right)-\cos \left(x\right)\::\begin{bmatrix}\mathrm{Solution:}\:&\:-\infty \:<x<\infty \\ \:\mathrm{Interval\:Notation:}&\:\left(-\infty \:,\:\infty \:\right)\end{bmatrix}"



"\mathrm{Range\:of\:}\sin \left(x\right)-\cos \left(x\right):\\\begin{bmatrix}\mathrm{Solution:}\:&\:-\sqrt{2}\le \:f\left(x\right)\le \sqrt{2}\:\\ \:\mathrm{Interval\:Notation:}&\:\left[-\sqrt{2},\:\sqrt{2}\right]\end{bmatrix}"



"\mathrm{Axis\:interception\:points\:of}\:\sin \left(x\right)-\cos \left(x\right):\quad \mathrm{X\:Intercepts}:\:\left(\frac{\pi }{4}+2\pi n,\:0\right),\:\left(\frac{5\pi }{4}+2\pi n,\:0\right),\:\mathrm{Y\:Intercepts}:\:\left(0,\:-1\right)"